Exercise. Use Abels theorem to find the general solultion. Note in each problem one solution is given. (a) t2y- 4ty +6y =
Answer #1

Guan dilleranded epnchan Spudt whansluan Pㅋ류 J- ce =CE ,4 =ct 4 Guen thut 9-67 integradng factv F) = e 1 General Doluthan epu+2 +9 =0 C4t piely+ley=o) PLE)=2 Bby Ables thesrem Spienlt Wlrovsleroan lC, ce -2dt -26 = ce -24- = Ce 2 ce 2 Ce nterady faGuen dolevandal euahan y(46+2) p lE7= -Spuedt WCY,, Ce .:By Alis Theorem ce RA- 42 4=26 -2logt loge PF= e = e Genaral Aoluha